# Dirac - von Neumann axioms in the setting of Continuous Model Theory

### v1

@@ 4 Hilbert space formalism and H-structures
CL-ultraproduct. We will work in a specific asymptotic class of H-structures of signature $(\mathcal{H}^{\mathrm{Def}},\mathrm{E})$ where the predicates of $\mathcal{H}^{\mathrm{Def}}$ on $\mathbb{U}(\mathfrak{n})$ are of the form
$\psi(k,\ldots,k_{m})=f(\frac{k_{1}}{\sqrt{\mathfrak{n}}},\ldots,\frac{k_{m}}{\sqrt{\mathfrak{n}}}),$
where $f(x_{1},\ldots,x_{m})$ is a smooth function ${\mathbb{R}}^{m}\to{\mathbb{C}}.$
The predicate with the same name $\psi$ on continuous $\mathbb{U}$ is
$\psi(x_{1},\ldots,x_{m}):=f(x_{1},\ldots,x_{m}).$
The quantifier $\mathrm{E}:=\{\mathrm{E}^{(m)}:m\in{\mathbb{N}}\},$ a family of quantifiers on finite intervals of diameter $2m<\sqrt{\mathfrak{n}},$ is defined as follows:
$\mathrm{E}^{(m)}_{k}\psi(k,\bar{p})=\mathrm{st}\left(\frac{1}{\sqrt{\mathfrak{n}}}\sum_{k\geq-m\sqrt{\mathfrak{n}}}^{k\leq m\sqrt{\mathfrak{n}}}\psi(k,\bar{p})\right),\ \mbox{for }\mathbb{U}(\mathfrak{n})$ (17)
$\mathrm{E}^{(m)}_{x}\psi(x,\bar{y})=\int_{-m}^{m}\psi(x,\bar{y})dx,\ \mbox{for }\mathbb{U}(\infty).$ (18)
@@
$\mathrm{E}^{\mathrm{loc}}_{k}\psi(k,\bar{p}):=\lim_{m\to\infty}\mathrm{E}^{(m)}_{k}\psi(k,\bar{p})$ (19)
Here, in case $\mathfrak{n}$ is finite, the limit should be understood as the value for the maximal $m$ satisfying $2m\leq\sqrt{\mathfrak{n}}.$
For a pseudo-finite $\mathfrak{n}$ it is assumed that $m$ runs in $\mathbb{N}$}, the standard positive integers. In the context of continuous model theory $\mathrm{E}^{\mathrm{loc}}_{k}$ is definable in terms of $\mathrm{E}^{(m)}_{k}$.
@@
4.7
Theorem. $(\mathcal{H},\mathrm{O})$ be a rigged Hilbert space with a family $\mathrm{O}$ of integral operators, and $\mathcal{H}^{\mathrm{Def}}$ its dense subspace closed under inner product and $\mathrm{O}.$
For any non-principal ultrafilter $\mathcal{D}$ on $\mathbb{N}$} the continuous model theory ultraproduct $(\mathbb{U}^{*},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ of finite H-structures $(\mathbb{U}(n),\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ is a continuous H-structure.
For every sentence $\sigma$ and every positive $\epsilon$ there is a subset $D_{\sigma,\epsilon}\in\mathcal{D}$ such that for all $n\in D_{\sigma,\epsilon}$ the value of $\sigma$ on $(\mathbb{U},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ differs from the value of $\sigma$ on $(\mathbb{U}(n),\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ by no more than $\epsilon.$
@@
$x:=\mathrm{st}(\frac{k}{\sqrt{\mathfrak{n}}}),$
(the standard part map). In particular, the interval $m\sqrt{\mathfrak{n}},m\sqrt{\mathfrak{n}}]$ in $\mathbb{U}(\mathfrak{n})$ corresponds to the interval [m,m][-m,m] in $\mathbb{R}$}.
@@
Next we prove the correspondence for quantifiers. It is enough to consider unary $\psi:{\mathbb{R}}\to{\mathbb{C}}.$ By definition $\psi(k)=f(\frac{k}{\sqrt{\mathfrak{n}}}),$ $f(x)$ smooth on $\mathbb{R}$}.
@@
$|\frac{1}{\sqrt{\mathfrak{n}}}\sum_{-m\sqrt{\mathfrak{n}}\leq k<m\sqrt{\mathfrak{n}}}f(\frac{k}{\sqrt{\mathfrak{n}}})\ -\ \int_{-m}^{m}f(x)dx|<\epsilon$
Indeed, the discrete formula is a Riemann sum with spacing $\Delta x=\frac{1}{\sqrt{\mathfrak{n}}}.$ By the left Riemann sums estimate for an interval $(a,b)$
@@
where $N$ is the number of points between $a=m$ and b=mb=-m and $M_{f}=\max\{f^{\prime}(x):b\leq x<a\}.$ Clearly, $M_{f}\frac{(2m)^{2}}{4m}\in{\mathbb{R}}$ and thus $M_{f}\frac{(2m)^{2}}{4m\sqrt{\mathfrak{n}}}<\epsilon$ because $\frac{1}{\sqrt{\mathfrak{n}}}$ is a non-standard infinitesimal.
@@ 5 Gaussian and perturbation-Gaussian states
$\psi(k_{1},\ldots,k_{m})=\eta\cdot\mathrm{e}^{-\pi i\frac{Q(k_{1},\ldots,k_{m})}{\mathfrak{n}}}=\eta\cdot\mathrm{e}^{-\pi iQ(\frac{k_{1}}{\sqrt{\mathfrak{n}}},\ldots,\frac{k_{m}}{\sqrt{\mathfrak{n}}})}$
@@
$\psi(x_{1},\ldots,x_{m})=\eta\cdot\mathrm{e}^{-\pi iQ(x_{1},\ldots,x_{m})},$
@@
$\psi(x,\bar{y})=\eta\cdot\mathrm{e}^{-\pi ic(\bar{y}))}\cdot\mathrm{e}^{-\pi i(ax^{2}+2xb(\bar{y})}$
@@
Lemma. For any d-dense subset $X\subset\mathbb{U}(\mathfrak{n})^{m}$ for any $\bar{y}\in{\mathbb{R}}^{m}$ there is $\bar{p}\in X$ such that $\mathrm{st}(\frac{1}{\sqrt{\mathfrak{n}}}\bar{p})=\bar{y}.$
Proof. When $p$ runs in $\mathbb{U}(\mathfrak{n})=[-\frac{\mathfrak{n}}{2},\frac{\mathfrak{n}}{2}]$ the numbers $\mathrm{st}(\frac{1}{\sqrt{\mathfrak{n}}}p)$ run continuously between -$\infty$ and $\infty$. Thus there is $\bar{p}\in\mathbb{U}(\mathfrak{n})^{m}$ such that $\mathrm{st}(\frac{1}{\sqrt{\mathfrak{n}}}\bar{p}).$
Density of $X$ implies that there is a tuple of non-negative standard integers $\bar{d}$ such that $\bar{p}+\bar{d}\in X.$ But $\mathrm{st}(\frac{1}{\sqrt{\mathfrak{n}}}\bar{d})=\bar{0}$ and thus we can assume $\bar{p}\in X.$ $\Box$
@@
Let $(\mathbb{U}^{*},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ be the ultraproduct constructed in 4.7, $\mathfrak{n}$ highly divisible, and $\mathrm{e}^{-\pi iax^{2}+2xb(\bar{y})}$ a Gaussian predicate on $\mathbb{U}^{*},$ where $a,b$ are rational. Then,
$\frac{1}{\sqrt{\mathfrak{n}}}\sum_{-\frac{\mathfrak{n}}{2a}<k\leq\frac{\mathfrak{n}}{2a}}\mathrm{e}^{-\pi i\frac{ak^{2}+2kb}{\mathfrak{n}}}=\int_{\mathbb{R}}\mathrm{e}^{-\pi i(ax^{2}+2xb)}dx$ (22)
@@
Proof. This is a direct consequence of 5.4 and 4.7. $\Box$
@@
Beyond free particles. Anharmonic oscillator. The Gaussian fragment of quantum mechanics modelled above (with a little more work includes also quantum harmonic oscillator) is the only part of QM which allows exact solutions. The more general version of QM would include states of the form $\mathrm{e}^{-i\frac{x^{2}+f(x)}{2\mathfrak{h}})}$ where $f(x)$ is a polynomial of degree >2.>2. In fact, the theory, due to physical and mathematical issues only deal with quite specific forms of such states. The key example is that of an anharmonic oscillator $\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathfrak{h}}}$ as analysed in [18]. This is also a much simplified analogue of so called $\phi^{4}$-quantum field theory.
@@
$\int_{\mathbb{R}}\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathfrak{h}}}dx,\ \ \ \lambda>0$
can only be carried out using perturbation methods, which impose specific restriction on coefficients, in particular $\mathfrak{h}$ have to be infinitesimally small in the example.
@@
$\mathrm{e}^{-\pi i\frac{H(k^{2}+\frac{1}{L}k^{4})}{\mathfrak{n}}}=\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathfrak{h}}},\ \ \lambda=\frac{\mathfrak{n}}{L}$
@@
$\mathrm{E}^{\mathrm{glob}}_{k}\psi=2\mathfrak{h}^{\frac{1}{2}}\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\mathrm{e}^{-\pi i\frac{Hk^{2}+\frac{H}{L}k^{4}}{\mathfrak{n}}}=$
$\mathfrak{h}^{\frac{1}{2}}\left(2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\mathrm{e}^{-\pi i\frac{Hk^{2}}{\mathfrak{n}}}+2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{H2}}\phi(k)\mathrm{e}^{-\pi i\frac{Hk^{2}}{2\mathfrak{n}}}\right)=$
$\mathfrak{h}^{\frac{1}{2}}(T_{0}(\mathfrak{h})+T_{\phi}(\mathfrak{h})).$
@@
$T_{0}(\mathfrak{h})=\mathrm{e}^{-\frac{\pi i}{4}}.$
So our aim is to evaluate $T_{\phi}(\mathfrak{h}).$
@@
Note that $k^{4}\leq(\frac{\mathfrak{n}}{H})^{4}$ and thus $\frac{Hk^{4}}{\mathfrak{n}^{2}}\leq 2\pi\mathfrak{h}(\frac{\mathfrak{n}}{H})^{2}$ and $\epsilon=o(\mathfrak{h}),$ so
$|\phi(k)|\leq\mathfrak{h}\cdot O(\lambda(\frac{\mathfrak{n}}{H})^{2})$
$T_{\phi}(\mathfrak{h})=|2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\phi(k)\mathrm{e}^{-\pi i\frac{Hk^{2}}{\mathfrak{n}}}|\leq|2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\phi(k)|\leq\mathfrak{h}\sqrt{\frac{H}{\mathfrak{n}}}\cdot\frac{\mathfrak{n}}{H}\cdot O(\lambda\frac{\mathfrak{n}^{2}}{H^{2}})=$
$\lambda\mathfrak{h}\cdot O(\frac{\mathfrak{n}}{H})^{2+1/2}$
@@
$\mathfrak{h}\to 0\mbox{ iff }\frac{\mathfrak{n}}{H}=O(1).$
@@
$T_{\phi}(H)\leq O(1)\cdot\lambda\mathfrak{h}\mbox{ when }\mathfrak{h}\to 0.$
@@
$\mathrm{E}^{\mathrm{glob}}_{k}\psi(k,\mathfrak{h})=\mathfrak{h}^{\frac{1}{2}}\sqrt{2\pi}\mathrm{e}^{\frac{\pi i}{4}}(1+I(\mathfrak{h})),\mbox{ where }I(\mathfrak{h})=O(1)\cdot\lambda\mathfrak{h}\mbox{ when }\mathfrak{h}\to 0.$
This is in a good agreement with $\mathrm{E}^{\mathrm{loc}}_{k}\psi(k,\mathfrak{h})$ calculated in [17] and [18] as an asymptotic (non-convergent) series of $\mathfrak{h}$.

### v2

@@ 4 Hilbert space formalism and H-structures
Other metric universes. Suppose ${\mathbb{M}}\subseteq{\mathbb{R}}^{m}=\mathbb{U}(\infty)^{m}$ is a real manifold definable in $(\mathbb{U}(\infty);\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ along with a metric $\mathrm{dist}_{M}$ and a measure $d^{\mathbb{M}}x$ on it, compatible with the metric and measure on $\mathbb{U}(\infty).$ Then bounded domains in $\mathbb{U}(\infty)^{m}$ induce bounded domains on ${\mathbb{M}}^{n}$ and we consider all the definable (in $(\mathbb{U}(\infty);\mathcal{H}^{\mathrm{Def}},\mathrm{O})$) continuos maps on bounded domains of ${\mathbb{M}}^{n}$ to be definable continuous predicates on domains in $\mathbb{M}$}.
Define $\Phi_{n}({\mathbb{M}})$ to be the space generated by all $\psi^{\prime}:{\mathbb{M}}^{n}\to{\mathbb{C}}$ where $\psi^{\prime}(x)=\psi(x)$ for a definable continuous predicate $\psi$ on a bounded domain $D_{\psi}\subseteq{\mathbb{M}}^{n},$ when $x\in D_{\psi}$, and $\psi^{\prime}(x)=0,$ when $x\notin D_{\psi}.$ The measure $d^{\mathbb{M}}x$ on $\mathbb{U}(\infty)$ induces the respective quantifier $\mathrm{E}^{\mathbb{M}}:\Phi_{n+1}({\mathbb{M}})\to\Phi_{n}({\mathbb{M}})$. As a result we have a well-defined inner product on $\Phi_{n}({\mathbb{M}})$ and the closure of $\Phi_{n}({\mathbb{M}})$ in the inner product topology gives us a Hilbert space $\mathcal{H}_{n}({\mathbb{M}})$ which is contained in the dual space $\Phi_{n}({\mathbb{M}})^{*}.$ This provides us with the rigged Hilbert space of quantum mechanics on the manifold $\mathbb{M}$}.
4.7
CL-ultraproduct. We will work with the asymptotic class of finite-fimensional $\mathcal{H}$-structures $\mathbb{U}(n)$ of signature $(\mathcal{H}^{\mathrm{Def}},\mathrm{O}).$
Note that the predicates of $\mathcal{H}^{\mathrm{Def}}$ on $\mathbb{U}(\mathfrak{n})$ can be written in the form
$\psi(k,\ldots,k_{m})=f(k_{1}\sqrt{\frac{2\pi}{\mathfrak{n}}},\ldots,k_{m}\sqrt{\frac{2\pi}{\mathfrak{n}}}),$
where $f(x_{1},\ldots,x_{m})$ is a differentiable function into $\mathbb{C}$} on finite diameter domains in ${\mathbb{R}}^{m}.$ (The predicate with the same name $\psi$ on the continuous $\mathbb{U}$ is just $f(x_{1},\ldots,x_{m})$.)
The quantifier $\mathrm{E}$ acting on generaised predicates $\psi$ can be represented as a family of quantifiers
$\mathrm{E}:=\{\mathrm{E}^{(m_{1},m_{2})}:m_{1}<m_{2}\in{\mathbb{Z}}\},$
on finite intervals of diameter $m_{2}-m_{1}<\sqrt{\frac{\mathfrak{n}}{2\pi}}$ is defined as follows:
$\mathrm{E}^{(m_{1},m_{2})}_{k}\psi(k,\bar{p})=\mathrm{st}\left(\sqrt{\frac{2\pi}{\mathfrak{n}}}\sum_{k\sqrt{\frac{2\pi}{\mathfrak{n}}}\geq m_{1}}^{k\sqrt{\frac{2\pi}{\mathfrak{n}}}\leq m_{2}}\psi(k,\bar{p})\right),\ \mbox{for }\mathbb{U}(\mathfrak{n})$ (23)
which corresponds to
$\mathrm{E}^{(m_{1},m_{2})}_{x}\psi(x,\bar{y})=\int_{m_{1}}^{m_{2}}\psi(x,\bar{y})dx,\ \mbox{for }\mathbb{U}(\infty).$ (24)
@@
$\mathrm{E}^{\mathrm{loc}}_{k}\psi(k,\bar{p}):=\lim_{m\to\infty}\mathrm{E}^{(-m,m)}_{k}\psi(k,\bar{p})$ (25)
when the limit exists.
Here, in case $\mathfrak{n}$ is finite, the limit should be understood as the value for the maximal $m$ satisfying $2m\leq\sqrt{\frac{\mathfrak{n}}{2\pi}}.$
For a pseudo-finite $\mathfrak{n}$ it is assumed that $m$ runs in $\mathbb{N}$}, the standard positive integers. In the context of continuous model theory $\mathrm{E}^{\mathrm{loc}}_{k}$ is definable in terms of $\mathrm{E}^{(-m,m)}$ which we will write as just $\mathrm{E}^{(m)}.$
@@
Recall that the definition of an $\mathcal{H}$-structure furnishes a definable constant $\mathbf{h}$ which in the finite-dimensional case can also be expressed in the form $\mathbf{h}(n)=\frac{h_{n}}{n},$ see (21).
4.8
Theorem. Let $\mathcal{D}$ be a non-principal ultrafilter on $\mathbb{N}$} such that
$\lim_{n/\mathcal{D}}\frac{h_{n}}{n}=\mathbf{h}(\infty).$
Then the continuous model theory ultraproduct $(\mathbb{U}^{*},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ of finite $\mathcal{H}$-structures $(\mathbb{U}(n),\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ is an $\mathcal{H}$-structure isomorphic to $(\mathbb{U}(\infty),\mathcal{H}^{\mathrm{Def}},\mathrm{O}).$
Equivalently, for every sentence $\sigma$ and every positive $\epsilon$ there is a subset $D_{\sigma,\epsilon}\in\mathcal{D}$ such that for all $n\in D_{\sigma,\epsilon}$ the value of $\sigma$ on $(\mathbb{U},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ differs from the value of $\sigma$ on $(\mathbb{U}(n),\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ by no more than $\epsilon.$
@@
$x:=\mathrm{st}(k\sqrt{\frac{2\pi}{\mathfrak{n}}}),$
(the standard part map). In particular, the interval $m\sqrt{\frac{\mathfrak{n}}{2\pi}},m\sqrt{\frac{\mathfrak{n}}{2\pi}}]$ in $\mathbb{U}(\mathfrak{n})$ corresponds to the interval [m,m][-m,m] in $\mathbb{R}$}.
@@
Next we prove the correspondence for quantifiers. It is enough to consider unary $\psi:{\mathbb{R}}\to{\mathbb{C}}.$ By definition $\psi(k)=f(k\sqrt{\frac{2\pi}{\mathfrak{n}}}),$ $f(x)$ differentiable on $\mathbb{R}$}.
@@
$|\sqrt{\frac{2\pi}{\mathfrak{n}}}\sum_{-m\leq k\sqrt{\frac{2\pi}{\mathfrak{n}}}<m}f(k\sqrt{\frac{2\pi}{\mathfrak{n}}})\ -\ \int_{-m}^{m}f(x)dx|<\epsilon$
Indeed, the discrete formula is a Riemann sum with spacing $\Delta x=\sqrt{\frac{2\pi}{\mathfrak{n}}}.$ By the left Riemann sums estimate for an interval $(a,b)$
@@
where $N$ is the number of points between $a=m$ and b=mb=-m and $M_{f}=\max\{f^{\prime}(x):b\leq x<a\}.$ Clearly, $M_{f}\frac{(2m)^{2}}{4m}\in{\mathbb{R}}$ and thus $M_{f}\frac{(2m)^{2}}{4m\sqrt{\mathfrak{n}}}<\epsilon$ because $\sqrt{\frac{2\pi}{\mathfrak{n}}}$ is a non-standard infinitesimal.
@@ 5 Gaussian and perturbation-Gaussian states
$\psi(k_{1},\ldots,k_{m})=\eta\cdot\mathrm{e}^{-\pi i\frac{Q(k_{1},\ldots,k_{m})}{\mathfrak{n}}}=\eta\cdot\mathrm{e}^{-i\frac{Q(k_{1}\sqrt{\frac{2\pi}{\mathfrak{n}}},\ldots,k_{m}\sqrt{\frac{2\pi}{\mathfrak{n}}})}{2}}$
@@
$\psi(x_{1},\ldots,x_{m})=\eta\cdot\mathrm{e}^{-i\frac{Q(x_{1},\ldots,x_{m})}{2}},$
@@
$\psi(x,\bar{y})=\eta\cdot\mathrm{e}^{-i\frac{c(\bar{y})}{2}}\cdot\mathrm{e}^{-i\frac{ax^{2}+2xb(\bar{y})}{2}}$
@@
Lemma. For any d-dense subset $X\subset\mathbb{U}(\mathfrak{n})^{m}$ for any $\bar{y}\in{\mathbb{R}}^{m}$ there is $\bar{p}\in X$ such that $\mathrm{st}(\sqrt{\frac{2\pi}{\mathfrak{n}}}\bar{p})=\bar{y}.$
Proof. When $p$ runs in $\mathbb{U}(\mathfrak{n})=[-\frac{\mathfrak{n}}{2},\frac{\mathfrak{n}}{2}]$ the numbers $\mathrm{st}(\sqrt{\frac{2\pi}{\mathfrak{n}}}p)$ run continuously between -$\infty$ and $\infty$. Thus there is $\bar{p}\in\mathbb{U}(\mathfrak{n})^{m}$ such that $\mathrm{st}(\sqrt{\frac{2\pi}{\mathfrak{n}}}\bar{p}).$
Density of $X$ implies that there is a tuple of non-negative standard integers $\bar{d}$ such that $\bar{p}+\bar{d}\in X.$ But $\mathrm{st}(\sqrt{\frac{2\pi}{\mathfrak{n}}}\bar{d})=\bar{0}$ and thus we can assume $\bar{p}\in X.$ $\Box$
@@
Let $(\mathbb{U}^{*},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ be the ultraproduct constructed in 4.8, $\mathfrak{n}$ highly divisible, and $\mathrm{e}^{-\pi iax^{2}+2xb(\bar{y})}$ a Gaussian predicate on $\mathbb{U}^{*},$ where $a,b$ are finite rational (possibly non-standard). Then,
$\sqrt{\frac{2\pi}{\mathfrak{n}}}\sum_{-\frac{\mathfrak{n}}{2a}<k\leq\frac{\mathfrak{n}}{2a}}\mathrm{e}^{-\pi i\frac{ak^{2}+2kb}{\mathfrak{n}}}=\int_{\mathbb{R}}\mathrm{e}^{-i\frac{ax^{2}+2xb}{2}}dx$ (28)
@@
Proof. This is a direct consequence of 5.4 and 4.8. $\Box$
@@
Beyond Gaussian. Anharmonic oscillator. The Gaussian fragment of quantum mechanics modelled above (with a little more work includes also quantum harmonic oscillator) is the only part of QM which allows exact solutions. The more general version of QM would include states of the form $\mathrm{e}^{-i\frac{x^{2}+f(x)}{2\mathbf{h}})}$ where $f(x)$ is a polynomial of degree >2.>2. In fact, the theory, due to physical and mathematical issues only deal with quite specific forms of such states. The key example is that of an anharmonic oscillator $\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathbf{h}}}$ as analysed in [18]. This is also a much simplified analogue of so called $\phi^{4}$-quantum field theory.
@@
$\int_{\mathbb{R}}\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathbf{h}}}dx,\ \ \ \lambda>0$
can only be carried out using perturbation methods, which impose specific restriction on coefficients, in particular $\mathbf{h}$ have to be infinitesimally small in the example.
@@
$\mathrm{e}^{-\pi i\frac{H(k^{2}+\frac{1}{L}k^{4})}{\mathfrak{n}}}=\mathrm{e}^{-i\frac{x^{2}+\lambda x^{4}}{2\mathbf{h}}},\ \ \lambda=\frac{\mathfrak{n}}{L}$
@@
$\mathrm{E}^{\mathrm{glob}}_{k}\psi=2\mathbf{h}^{\frac{1}{2}}\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\mathrm{e}^{-\pi i\frac{Hk^{2}+\frac{H}{L}k^{4}}{\mathfrak{n}}}=$
$\mathbf{h}^{\frac{1}{2}}\left(2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\mathrm{e}^{-\pi i\frac{Hk^{2}}{\mathfrak{n}}}+2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{H2}}\phi(k)\mathrm{e}^{-\pi i\frac{Hk^{2}}{2\mathfrak{n}}}\right)=$
$\mathbf{h}^{\frac{1}{2}}(T_{0}(\mathbf{h})+T_{\phi}(\mathbf{h})).$
@@
$T_{0}(\mathbf{h})=\mathrm{e}^{-\frac{\pi i}{4}}.$
So our aim is to evaluate $T_{\phi}(\mathbf{h}).$
@@
Note that $k^{4}\leq(\frac{\mathfrak{n}}{H})^{4}$ and thus $\frac{Hk^{4}}{\mathfrak{n}^{2}}\leq 2\pi\mathbf{h}(\frac{\mathfrak{n}}{H})^{2}$ and $\epsilon=o(\mathbf{h}),$ so
$|\phi(k)|\leq\mathbf{h}\cdot O(\lambda(\frac{\mathfrak{n}}{H})^{2})$
$T_{\phi}(\mathbf{h})=|2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\phi(k)\mathrm{e}^{-\pi i\frac{Hk^{2}}{\mathfrak{n}}}|\leq|2\sqrt{\frac{H}{\mathfrak{n}}}\sum_{0\leq k<\frac{\mathfrak{n}}{2H}}\phi(k)|\leq\mathbf{h}\sqrt{\frac{H}{\mathfrak{n}}}\cdot\frac{\mathfrak{n}}{H}\cdot O(\lambda\frac{\mathfrak{n}^{2}}{H^{2}})=$
$\lambda\mathbf{h}\cdot O(\frac{\mathfrak{n}}{H})^{2+1/2}$
@@
$\mathbf{h}\to 0\mbox{ iff }\frac{\mathfrak{n}}{H}=O(1).$
@@
$T_{\phi}(H)\leq O(1)\cdot\lambda\mathbf{h}\mbox{ when }\mathbf{h}\to 0.$
@@
$\mathrm{E}^{\mathrm{glob}}_{k}\psi(k,\mathbf{h})=\mathbf{h}^{\frac{1}{2}}\sqrt{2\pi}\mathrm{e}^{\frac{\pi i}{4}}(1+I(\mathbf{h})),\mbox{ where }I(\mathbf{h})=O(1)\cdot\lambda\mathbf{h}\mbox{ when }\mathbf{h}\to 0.$
This is in a good agreement with $\mathrm{E}^{\mathrm{loc}}_{k}\psi(k,\mathbf{h})$ calculated in [17] and [18] as an asymptotic (non-convergent) series of $\mathbf{h}$.